OpenStudy (anonymous):

Don has been tracking the population of a town. The population was 7,220 last year and has been increasing by 1.45 times each year. How can Don develop a function to model this

OpenStudy (anonymous):

A) f(x) = (7220 • 1.45)x where 1.45 is the rate of growth B) f(x) = 1.45(7220)x where 7220 is the rate of growth C) f(x) = 7220(1.45)x where 1.45 is the rate of growth D)f(x) = 7220(x)1.45 where 7220 is the rate of growth

OpenStudy (anonymous):

The answer should be: \[f(x) = 7220 \times (1.45)^x\] Where 1.45 is the growth rate.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

You're welcome!! Xo!